OpenStudy (anonymous):

Please help!!!!!!! I will reward you with e medal just please help!!!!!!!!!!!!

OpenStudy (anonymous):

Solve using two different methods. Explain which method you found to be more efficient. a. 3x – 9y = 3 b. 7x – 3y = 20 c. y = 1/2x - 6 6x – 3y = -24 5x + 3y = 16 2x + 6y = 19

OpenStudy (anonymous):

I would help you with this but I took this in middle school! and then my freshman year and now im a senior. I forgot. I'm so sorry

OpenStudy (anonymous):

Someone will help you though!

OpenStudy (anonymous):

=( its ok i just dont know and i want it done but i cant figure it out! But thanks anyway!

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

YAY Please help I need help with b and c...

OpenStudy (anonymous):

have you taken the first equation and isolated one variable yet?

OpenStudy (anonymous):

Substitution method 3x - 9y = 3 Dividing throughout by 3,we get x - 3y = 1 x = 1 + 3y ..... ( i ) 6x - 3y = - 24 .....( ii ) that's all i know

OpenStudy (anonymous):

i have no clue what that means @swmchick

OpenStudy (anonymous):

actually ignore that

OpenStudy (anonymous):

if you look at the two problems, you notice one has a -3y and the other is +3y right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so would i add both equations together then

OpenStudy (anonymous):

so what you can do is set up one giant addition problem that will look like this \[(7x-3y=20)-5x=3y=16\]

OpenStudy (anonymous):

exactly. just add all like variables and its easy to solve for x

OpenStudy (anonymous):

ok i got x now what do i do to get y?

OpenStudy (anonymous):

then all you have to do is plug that value back into any equation to solve for y

OpenStudy (anonymous):

now is that substitution?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

tell me when you got it and we can move to c

OpenStudy (anonymous):

ok so for x i got x=3 and for y i got y = -1 2/3 is that correct?

OpenStudy (anonymous):

sounds right

OpenStudy (anonymous):

to check just plug it into the other equation :)

OpenStudy (anonymous):

ok now for c would we do elimination?

OpenStudy (anonymous):

Because I ave no clue how to do that lol.

OpenStudy (anonymous):

you would do the same thing you did for b. it just requires one more step. you would rearrange the first equation to 1/2x - y = 6 Then multiply that whole equation by -4

OpenStudy (anonymous):

you would get -2x=4y=-26

OpenStudy (anonymous):

then you can make both of the equations into one big addition problem and solve the same way you did for b

OpenStudy (anonymous):

you mean -2x + 4y = -24

OpenStudy (anonymous):

yes. sorry. typo

OpenStudy (anonymous):

are you all good? :)

OpenStudy (anonymous):

ok i am solving right now give me just a second.

OpenStudy (anonymous):

can you solve my problem after?

OpenStudy (anonymous):

i can try yes.

OpenStudy (anonymous):

ok so i got x=11 and y= -1/2 is that right and did we use substitution again?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

KEWL! Thanks ok so give me the link to yours and ill see if i can solve it.